OpenStudy (anonymous):

F(x)=x+6/3 and g(x)=1x-2 Part 2. Show your work to prove that the inverse of f(x) is g(x). How they are inverses is because they have to to undo each other and g(f(x))=x G(f(x))=g(x+6/3)=1(x+6/3)=x+2-2=x and F(g(x)=f(1x-2)=(x-2)+6/3 because x-2+2=x Part 3. Show your work to evaluate g(f(x)). (SHOWED IN PART 2) g(f(x))=x G(f(x))=g(x+6/3)=1(x+6/3)=x+2-2=x and F(g(x)=f(1x-2)=(x-2)+6/3 Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph

OpenStudy (anonymous):

@ganeshie8 you said you would help all i need help is with part 4!!

OpenStudy (anonymous):

@campbell_st i need help with part 4 if you can

OpenStudy (anonymous):

@myininaya i dont get part 4 the graphing and making the table

myininaya (myininaya):

it sounds like it wants you to make a table like x|f(x)|g(x) --------- 0| 1| 2| 3| 4|

myininaya (myininaya):

plug in 0 into both functions and see what the output is

myininaya (myininaya):

then 1

myininaya (myininaya):

then 2

myininaya (myininaya):

and so on...

myininaya (myininaya):

that is how I interpret it

OpenStudy (anonymous):

okay so what i dont get is what to graph and how to

myininaya (myininaya):

you graph both f(x) and g(x)

myininaya (myininaya):

and also y=x

myininaya (myininaya):

so in all you will graph 3 different functions on the same little picture thing

myininaya (myininaya):

\[f(x)=x+\frac{6}{3} \text{ or } f(x)=\frac{x+6}{3} ?\]

OpenStudy (anonymous):

but there is only 2 different functions i thought?

OpenStudy (anonymous):

and the first one

myininaya (myininaya):

But it wants to graph three f(x),g(x), and finally y=x

OpenStudy (anonymous):

ohh okay so how do i do that?

myininaya (myininaya):

that is three functions in all

myininaya (myininaya):

so f(x)=x+2

myininaya (myininaya):

Graph f(x)=x+2

OpenStudy (anonymous):

how and idk where to graph lol im in virutal is there a site i can use?

myininaya (myininaya):

to graph?

myininaya (myininaya):

use paper and pencil

OpenStudy (anonymous):

and i need help on the table i made it with all the numbers plugged in and yes i know how to graph but i dont know what site to use

myininaya (myininaya):

f(x)=mx+b is a linear function with slope m and y-intercept b

OpenStudy (anonymous):

so the slope is 1 and y intercept is 2

myininaya (myininaya):

yes for f(x)=x+2

OpenStudy (anonymous):

first can we do the table because when i plugged in 1 i got 2.3 repeating

myininaya (myininaya):

for f?

OpenStudy (anonymous):

yes

myininaya (myininaya):

if x=1 then f(1)=1+2=3 and g(1)=1-2=-1

myininaya (myininaya):

and the last function y=x we can call that h(x)=x h(1)=1

OpenStudy (anonymous):

no it would be f(1)=1+6/3 i thought then add 1 to 6 and then divide 7 by 3?

myininaya (myininaya):